OpenStudy (anonymous):

Again substitution method! 2x+9y=66 -6x+y=26 what is the order pair?

OpenStudy (anonymous):

Solve for one variable and then plug that in to the other equation to get an equation with one unknown.

OpenStudy (anonymous):

ok can you please help me more with that

OpenStudy (anonymous):

Ok so I would take -6x+y=26 and solve for y and then plug that into 2x+9y=66 for y and solve for x.

OpenStudy (anonymous):

Well that being said I got y=6x=26 for the y= and for the x= I got y/6-13/3 is that correct cause I am not positive

OpenStudy (anonymous):

Not quite sure what you did, but I'll start working it out. Maybe you can see it that way. :)

OpenStudy (anonymous):

y=6x+26 so then the other equation becomes 2x+9(6x+26)=66

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

Can you solve for x from there?

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

Yahtzee! Now plug in -3 for x into let's say y=6x+26 to get the y value. Then your ordered pair is (x,y)

OpenStudy (anonymous):

8

OpenStudy (anonymous):

-3,8 is the order pair

OpenStudy (anonymous):

So how do you know which equation to put in so that you get the right answer or does it work with both?

OpenStudy (anonymous):

It will work with both.

OpenStudy (anonymous):

I hope that is right I gave you a medal for that thanks a bunch I am trying to learn this math and it is not easy to me but I am welling to set here til I can work it out with someone

OpenStudy (anonymous):

That was correct thanks again!!!!!

OpenStudy (anonymous):

Thanks. It is right :) Solving systems of equations by substitution method is the same process every time. No problem.